For the reaction 2H2 + O2 2H2O, what is the total mass in grams of water formed when 8.00 grams of hydrogen reacts completely with excess oxygen?2
1
Expert's answer
2020-04-06T11:49:56-0400

Let's start by writing the balanced reaction equation:

2H2 + O2 "\rightarrow" 2H2O

As one can see, when there is enough oxygen, 2 molecules of hydrogen give 2 molecules of water. Therefore, when there is enough oxygen, the number of the moles of hydrogen reacted equals the number of the moles of water produced:

"n \text{ (H}_2\text{)} = n \text{ (H}_2\text{O)}"

Let's calculate the number of the moles of hydrogen in 8.00 grams of hydrogen (the number of the moles is mass "m" divided by molar mass "M", "M"(H2) =2.01568 g/mol):

"n \text{ (H}_2\text{)}= \frac{m}{M} = \frac{8.00}{2.01568} = 3.97" mol

"n \text{ (H}_2\text{O)} = 3.97" mol

The molar mass of water is 18.01528 g/mol. The mass of water produced is:

"m = n\cdot M = 3.97 \cdot 18.01528 = 71.5" g

Answer: The total mass in grams of water formed when 8.00 grams of hydrogen reacts completely with excess oxygen is 71.5 g.